OpenStudy (anonymous):

Christian is comparing two checking accounts. One has a monthly fee of $7 and a per-check fee of $0.20, and the other has a monthly fee of $8 and a per-check fee of $0.15. What is the minimum number of checks Christian needs to write for the second bank to be a better option? A. 19 B. 21 C. 18 D. 22 @Loveiskey18 @jeherohaku @InYourHead @HomeschoolGrad @ganeshie8

OpenStudy (anonymous):

i think its b @Loveiskey18

OpenStudy (anonymous):

How did you come to that conclusion?

OpenStudy (anonymous):

lol idk i quessed

OpenStudy (anonymous):

@Loveiskey18 u there?

OpenStudy (anonymous):

7 + .20x = 8 + .15x .20x - .15x = 8 - 7 .05x = 1 x = 1/.05 x = 20 They are equal when writing 20 checks. Therefore, when writing 19 checks, A is better. check... 7 + .20x when writing 19 checks 7 + .20(19) 7 + 3.8 = 10.80 <-- this is better because it is cheaper 8 + .15x when writing 19 checks 8 + .15(19) 8 + 2.85 = 10.85

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

no thank you :)

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

this is also the wrong answer, the correct answer is 21 not 20